OpenStudy (anonymous):

solve square root x + 4 = -10

OpenStudy (aravindg):

Square on both sides.

OpenStudy (anonymous):

You can't take the square of a negative number... otherwise you'll come up with an imaginary number.

OpenStudy (anonymous):

You can square a negative number. \((-10)^2=(-10)(-10)\)

OpenStudy (anonymous):

What I mean is the square root of -10.

OpenStudy (aravindg):

?

OpenStudy (anonymous):

You square both sides to get rid of the square root.

OpenStudy (anonymous):

But either way, it does look like there is no solution.

OpenStudy (anonymous):

Wolfram Alpha responds with no solution. See the attachment.

OpenStudy (kainui):

square root x + 4 = -10 means either\[\sqrt{x}+4=-10\] or \[\sqrt{x+4}=-10\]

OpenStudy (kainui):

So x is either 14^2 or 96 specifically.